OpenStudy (anonymous):

For the function y= -1 + 6 cos( (2 pi/7)(x-5) ), what is the maximum value? For the function y = -2 + 5 sin(pi/12(x-2) ), what is the minimum value?

OpenStudy (anonymous):

the maximum of cosine is 1 the maximum of 6 times cosine is 6 the maximum of 6 times cosine - 1 is ...

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

\[\left| \cos \theta \right|\le1,\] max. value of y=-1+6*1=5

OpenStudy (anonymous):

min.value of sine=-1 y=-2-5=-7